Find the equation of the normal line to the curve of the equation x 2 y + xy2 = 6 at the point (2, 1).


1
Expert's answer
2021-10-26T11:57:40-0400
"x^2y+xy^2=6"

Differentiate both sides with respect to "x"


"\dfrac{d}{dx}(x^2y+xy^2)=\dfrac{d}{dx}(6)"

Use the Chain Rule


"2xy+x^2 \dfrac{dy}{dx}+y^2+2xy\dfrac{dy}{dx}=0"

Solve for "\dfrac{dy}{dx}"


"\dfrac{dy}{dx}=-\dfrac{2xy+y^2}{x^2+2xy}"

Find the slope of the tangent line to the curve at the point "(2, 1)"


"{slope}_1=m_1=-\dfrac{2(2)(1)+(1)^2}{(2)^2+2(2)(1)}=-\dfrac{5}{8}"

Find the slope of the normal line to the curve at the point "(2, 1)"


"{slope}_2=m_2=-\dfrac{1}{m_1}=\dfrac{8}{5}"

The equation of the normal line in point-slope form


"y-1=\dfrac{8}{5}(x-2)"

The equation of the normal line  to the curve of the equation "x ^2 y + xy^2 = 6" at the point "(2, 1)" in slope-intercept form


"y=\dfrac{8}{5}x-\dfrac{11}{5}"